# Efficacy of Image Similarity as a Metric for Augmenting Small Dataset Retinal Image Segmentation

### v1

@@ Appendix A Generated Image Examples
Fig. A.1, A.2, and A.3 show example synthetic images which have been randomly selected from datasets with FIDs of 256.76, 81.47, and 57.50 respectively.
@@ Appendix B Section VI-B Best-fit Parameters
Below is a table of the best-fit parameters for the curves shown in Section 6.2. Parameters were bound to lie between the priors shown in (6).
@@
Table B.1: Best-fit parameters for the best-fit curves presented in Fig. 6.
@@ Appendix C Carbon Impact of Experiments
Machine Learning requires continuous use of energy intensive computing hardware. Here we wish to highlight a rough estimate of the carbon dioxide equivalent emissions purely from running the experiments described above. Estimates were carried out by using the python package CodeCarbon (Courty et al., 2024). Monitoring a single experiment during Section 5 we found a carbon emission rate of 0.0127 gCO2e/s. We assume that, since we are performing effectively the same experiment repeatedly, that this emission rate will be the same for all.

### v3

@@ Appendix A Generated Image Examples
Fig.˜A.1, A.2, and A.3 show example synthetic images which have been randomly selected from datasets with FIDs of 256.76, 81.47, and 57.50 respectively.
@@ Appendix B Section VI-B Best-fit Parameters
Below is a table of the best-fit parameters for the curves shown in Section˜6.2. Parameters were bound to lie between the priors shown in (6).
@@
Table B.1: Best-fit parameters for the best-fit curves presented in Fig.˜6.
@@ Appendix C Carbon Impact of Experiments
Machine Learning requires continuous use of energy intensive computing hardware. Here we wish to highlight a rough estimate of the carbon dioxide equivalent emissions purely from running the experiments described above. Estimates were carried out by using the python package CodeCarbon (Courty et al., 2024). Monitoring a single experiment during Section˜5 we found a carbon emission rate of 0.0127 gCO2e/s. We assume that, since we are performing effectively the same experiment repeatedly, that this emission rate will be the same for all.